OpenStudy (anonymous):

Change 3.2513513513513513513513... to a fraction in simplest form. Please show me how to do it.

jimthompson5910 (jim_thompson5910):

Let x = 3.25135135135135... Then 100x = 325.135135135135... and 100000x = 325135.135135135... Now subtract the two 100000x - 100x = 325135.135135135... - 325.135135135135... 99900x = 324810 and I'm pretty sure you know how to do the rest...

OpenStudy (anonymous):

1203/370

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Converting a recurring decimal to vulgar fraction A decimal with recurring value is called recurring decimal. E.g: 2/9 will give 0.22222222...... where 2 is recurring number. Method: a) Separate the recurring number from the decimal fraction. b) Annex denominator with '9' as many times as the length of the recurring number. c) Reduce the fraction to its lowest terms. Example: Consider 0.2323232323 Step a: The recurring number is 23 Step b: 23/99 [the number 23 is of length 2 so we have added two nines] Step c: Reducing it to lowest terms : 23/99 [it can not be reduced further]. How to Convert a mixed-recurring decimal to vulgar fraction ? A decimal with both recurring and non-recurring value is called mixed recurring decimal. E.g: 28/25 will give 1.1199999999...... where 11 is non-recurring number and 9 is recurring number. Method: a) Separate the recurring number, non recurring number from the decimal fraction. b) Round the decimal after point to the first recurring value. c) Result of step b - non recurring number. d) Annex as many '0' as non-recurring number length and as many '9' as recurring number length. e) Step c / Step d f) Add the fraction with the number before decimal point. Example: Consider 1.11999999... Step a: The recurring number is 11, non-recurring number is 9 Step b and c: 119-11 [rounded value of number after decimal point - non recurring value] Step d: 900 Step e: 108/900 [c/d] Step f: 1+108/900 [adding with number before decimal point ] Reducing it to lowest terms : 900+108 / 900 = 1008/900 = 28/25.

OpenStudy (anonymous):

Oh ok. Do you get to 100000 before you subtract the two decimal places because that's where the infinite or recurring number starts?

jimthompson5910 (jim_thompson5910):

Yes, you locate when the terms start repeating again. That's why I moved the decimals in the positions that I did.

jimthompson5910 (jim_thompson5910):

And moving the decimal to the right is the same as multiplying by 10.

OpenStudy (anonymous):

Alright. Thank you! :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Alright. Thank you! :)

OpenStudy (anonymous):

Alright. Thank you! :)